Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-page---20131025

Observers: Massimo Turatto, Susanna Spiro
Support team: S.Benetti, N.Elias-Rosa
TAT on duty: Morgan Fraser, Kate Maguire
------------------------------------------------

Subrun 1, night 3

Instruments: SOFI

WEATHER CONDITION
beginning: wind speed 9 m/s; humidity 55 % ; despejado- photometric
middle: wind speed  9 m/s; humidity 22 %;  clear
end: wind speed  10 m/s; humidity 14 %;  clear

afternoon CALIBRATIONS
PESSTO_SOFI_specflat
PESSTO_SOFI_FlatJHK              twice

OBSERVATIONS: start regularly at sunset.
Target                       Type             UT         Filter/Grism      Slit            Seeing    Grade (A/B/C)     Comments

Hip094154              tellSTD         23:36           BG,RG           1.0             1.5              A
SN2012ca              follow-up      00:15            RG                1.0            1.6-1.9        B                 stop at 1.20h exptime, seeing too bad
Hip097557             tellSTD         02:30           BG,RG           1.0              2.0              A
SN2013fs               follow-up      02:44           JHKs                                1.9              A
lost  1.5 h for technical problems. NTT got stuck.
PSNJ01591443+1901339  follow-up   04:30   JHKs                                 1.5-1.7       A
Ogle-2013-SN-079   follow-up    06:30          J                                      1.8             A                  SN barely visible, if any
SN2013ej                follow-up       07:30          BG,RG           1.0             1.5              A                 too low for getting JHKs
Hip016879             tellSTD            07:50          BG,RG           1.0             1.5              A
SN2013fc                follow-up       08:10          JHKs                                1.4             A                   
Ogle-2013-SN-079   follow-up    08:30          J x 2                                1.5             A                  repeated with improved conditions
LTT3218                 fluxSTD         08:50          BG,RG            1.0                               A

morning CALIBRATIONS
PESSTO_SOFI_SpecArc    





Comments